OpenStudy (anonymous):

Which of the following verbal expressions represents the algebraic expression 2x – 8? A. eight less than twice a number B. the difference between eight and twice a number C.twice eight less than a number D.twice a number and eight

OpenStudy (anonymous):

D=2(x+8) C=2(8)-x B=8-2x

OpenStudy (anonymous):

thus the answer is…

OpenStudy (anonymous):

C

OpenStudy (anonymous):

no

OpenStudy (anonymous):

B

OpenStudy (anonymous):

D,C,B are wrong as i showed above^

OpenStudy (anonymous):

A

OpenStudy (anonymous):

A would b the correct answer now may i have a medal

OpenStudy (anonymous):

no click the best response button ------------------------------->